OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 78. (use 3.14 for pi)

OpenStudy (anonymous):

s = 3.14(4)(39)

OpenStudy (anonymous):

surface area of sphere = 4*3.14* r^2 r = radius = 78/2 = 39

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

s = 3.14(4)(39)